Diane invests $3500 part at 3%, part at 4%. How much does she invest at each rate if she receive in one year interest at $115?
Answer by lwsshak3(11628)

Diane invests $3500 part at 3%, part at 4%. How much does she invest at each rate if she receive in one year interest at $115?
***
let x=amt invested at 3%
3500-x=amt invested at 4%
3%x+4%(3500-x)=115
.03x+.04(3500-x)=115
.03x+140-.04x=115
.01x=25
x=2500
3500-x=1000
amt invested at 3%=$2500
amt invested at 4%=$1000